I have 4 oz. of liquid which contains a 60 % hydrogen peroxide solution. How much water should I add to reduce the solution to a 3 % hydrogen peroxide solution?
Answer by Alan3354(69443)

I have 4 oz. of liquid which contains a 60 % hydrogen peroxide solution. How much water should I add to reduce the solution to a 3 % hydrogen peroxide solution?
------------
4 oz of 60% = 4*0.6 = 2.4 oz of H2O2
----
2.4 oz = 3% of the solution
Total solution = 2.4/0.03 = 80 oz
Add 76 oz of water
----------------
Or,
4*60 + w*0 = (4+w)*3
240 = 3w + 12
3w = 228
w = 76 oz